36. Simplify the sums and differences. Give exact answers. (3sqrt2)(-4sqrt10)
88. Simplify each expression. sqrt2t^5 *sqrt10t^4

Answer by stanbon(75887)

36. Simplify the sums and differences. Give exact answers.
(3sqrt2)(-4sqrt10)
---
= -12 sqrt(20)
= -12sqrt(4)*sqrt(5)
= -12*2*sqrt(5)
= -24sqrt(5)
=================

88. Simplify each expression.
sqrt(2t^5) *sqrt(10t^4)
---
= sqrt(20t^9)
---
= sqrt(4t^8)*sqrt(5t)
---
= 2t^4*sqrt(5t)
======================
Cheers,
Stan H.